July 23, 1976
Work continued on cut #3, grids N-O 40-42, on the yellow earth stratum. Work also continued on grids I 38-42, and on the middle room. In the middle room grids O-P 32-37 were cleaned. The concentration of coarseware in grids JK 31-35 was covered with plastic sheeting and the trench was refilled with earth. The area is being left for another season, when there is time for excavation and appropriate storeroom space. Two boxes of pottery taken uot of this area will be boxed and stored.
- Find #1
- I 40
- red stratum
I casette of tile dumped.
There is no page 165
9 AM- work discontinued because of rain.
- Find #2
- O 41
- yellow soil
Section of worked antler
|Property or Relation||Value(s)|
[Standard: Dublin Core Terms]
Open Context editors work with data contributors to annotate datasets to shared vocabularies, ontologies, and other standards using 'Linked Open Data' (LOD) methods.
The annotations presented above approximate some of the meaning in this contributed data record to concepts defined in shared standards. These annotations are provided to help make datasets easier to understand and use with other datasets.
To the extent to which copyright applies, this content
carries the above license. Follow the link to understand specific permissions
Required Attribution: Citation and reference of URIs (hyperlinks)